### v2

# Route to Chaos in Multi-Species Ecosystems

Abstract
We investigate species-rich mathematical models of ecosystems. While much of the existing literature focuses on the properties of equilibrium fixed points, persistent dynamics (e.g., limit cycles or chaos) have also been observed, both in natural or lab-controlled ecosystems and in mathematical models. Here we emphasize the emergence of limit cycles following Hopf bifurcations tuned by the variability of interspecies interaction. As this variability increases, and owing to the large dimensionality of the system, limit cycles typically acquire a growing spectrum of frequencies. This often leads to the appearance of strange attractors, with a chaotic dynamics of species abundances characterized by a positive Lyapunov exponent. We observe that limit cycles and strange attractors preserve biodiversity to some extent, as they maintain dynamical stability without species extinction. We give numerical evidences that this route to chaos dominates in ecosystems with strong enough interactions and where predator-prey behavior dominates over competition and mutualism. Based on arguments from random matrix theory, we further conjecture that this scenario is generic in ecosystems with large number of species, and identify the key parameters driving it. Overall, we show that the model we consider provides a unifying framework, where a wide range of population dynamics emerge from a simple few-parameter model.
The time evolution of a species population in a given ecosystem can follow very diverse paths. Of course, a population can remain relatively stable over time, but it can also oscillate between highs and lows, from one year to another. More surprisingly, recent studies have shown that this evolution can even be chaotic – meaning that no regular patterns emerge, and the population at any given time is highly sensitive to its initial state. We show that the three behaviors aforementioned (steady state, cycle, and chaos) can be modeled in a parcimonious version of the generalized Lotka-Volterra model. Furthermore, we identify the two model parameters that govern the transition between these three states. Namely, we observe that, as the interaction strengths between species increases, ecosystems tend to transition from a steady state, to a cyclic trajectory, and further to a chaotic behavior.
I Introduction
One of the main challenges in theoretical ecology is to connect predictions from mathematical models of population dynamics to empirical observations of species coexistence in natural or laboratory-controlled ecosystems. May, Robert and McLean, Angela R (2007) In the latter it is established that individual populations fluctuate in time, often with large qualitative and quantitative differences between species. Ranta, Kaitala, and Lundberg (1998); Lundberg et al. (2000); Inchausti and Halley (2003); Vandermeer (2006) In some instances, population abundances exhibit synchronized, periodic oscillations, while in others population dynamics appears chaotic. Hastings et al. (1993); Ellner and Turchin (1995); Hunter and Price (1998); Huisman and Weissing (1999); Benincà et al. (2008); Rogers, Johnson, and Munch (2022); Prendergast et al. (2010); Pearce, Agarwala, and Fisher (2020) It is highly desirable to find out whether the wide range of observed dynamics can be captured in a single mathematical model, by varying few parameters, and if yes, what are the key characteristics this model should retain. Investigating mathematical models can furthermore shed light on fundamental qualitative questions such as whether populations fluctuations are endogenous or exogenous, i.e., if they are generated by intrinsic interactions or by external sources. Ellner and Turchin (1995); Scheffer et al. (2003)
In this manuscript, we numerically investigate large Lotka-Volterra models with random interactions, which are standard multi-species models of population dynamics. We emphasize the richness of their dynamics as a function of two key parameters which are (i) the variability $\sigma$ of interspecies interactions, and (ii) the cross-diagonal covariance parameter $\gamma$ of the interaction matrix. Our main finding is that, for a predominance of predator-prey pairs of species, the stable fixed-points prevailing at weak interaction variability generically lose their stability through Hopf bifurcations. Limit cycles emerge, where surviving species have periodically oscillating abundances. At still stronger interactions, strange attractors appear, possibly from cascades of bifurcations, which lead to a chaotic dynamics of population abundances characterized by a positive largest Lyapunov exponent. This route to chaos illustrates how stationarity, oscillating periodicity and chaos in the dynamics of species abundances exist in rather general models of population dynamics, depending on $\sigma$ and $\gamma$. One important result is that all observed population dynamics in multi-species ecosystems can be reproduced by a unified and parcimonious mathematical model.
Much of the existing literature on theoretical ecology is based on the surmise that the observed states and dynamics of ecosystems can be described by the time-asymptotic behavior of mathematical models. The latter is characterized by either a time-independent, stationary distribution of population abundances, or by persistently varying abundances, oscillating periodically or aperiodically. Accordingly, there has been a significant focus on the equilibrium fixed points of large systems of competing species, their stability and feasibility. May (1972); Roberts (1974); Rieger (1989); Allesina and Tang (2012); Tang, Pawar, and Allesina (2014); Fyodorov and Khoruzhenko (2016); Clenet, Massol, and Najim (2023); Ros et al. (2023) One key parameter is the variability $\sigma$ of the interspecies interaction. It has been found that a unique, asymptotically stable fixed point exists at weak enough $\sigma<\sigma_{c}$May (1972); Rieger (1989); Clenet, Massol, and Najim (2023); Ros et al. (2023) and that $\sigma_{c}$ increases in ecosystems with dominating predator-prey interactions. Allesina and Tang (2012); Tang, Pawar, and Allesina (2014) For $\sigma>\sigma_{c}$, one enters a phase with multiple unstable equilibria. Ros et al. (2023) Stabilization may still occur via species extinction, which drives the ecosystem to a novel, stable fixed point with reduced biodiversity. Beyond fixed points, the emergence of limit cycles through a Hopf bifurcation has been emphasized, albeit for small systems with only few species. Gilpin (1975); Coste, Peyraud, and Coullet (1979); Gardini, Lupini, and Messia (1989); Fussmann et al. (2000); Mcgehee et al. (2008) Further regimes with aperiodic persistent dynamics have recently been highlighted at stronger interactions and/or larger number of species, Hu et al. (2022) in ecosystems with sparse interactions, Roy et al. (2020); Marcus, Turner, and Bunin (2024) or communities with migrations. Dalmedigos and Bunin (2020)
In contrast to this mainstream philosophy focusing on the time-asymptotic behavior of mathematical models, several works have advocated that sufficiently large ecosystems have long relaxation times so that, by nature, they are always observed in a transient state. Hastings and Higgins (1994); Schreiber (2003); Morozov, Banerjee, and Petrovskii (2016); Hastings et al. (2018) Such large relaxation times may follow for instance from the slowing down of the dynamics in ecosystems close to criticality. Solé et al. (2002); Mora and Bialek (2011) Below we show that all these different behaviors naturally emerge in different regimes of a single mathematical model, as a function of only two parameters.
II Model and method
Population dynamics in multi-species ecosystems is commonly studied in the framework of the generalized Lotka-Volterra model. Goel, Maitra, and Montroll (1971) Here we consider the variant investigated in Ref. Ros et al., 2023, which reads
$\dot{N}_{i}=N_{i}\left[\kappa_{i}-N_{i}-\frac{\mu}{S}\sum_{j=1}^{S}N_{j}-\frac{\sigma}{\sqrt{S}}\sum_{j=1}^{S}{\mathbb{A}}_{ij}N_{j}\right]\,.$ (1)
Equation (1) determines the time-evolution of the normalized abundance $N_{i}(t)\geq 0$ of species $i=1,2,\ldots S$, as a function of its intrinsic growth rate $\kappa_{i}$ and its interaction with other species. Interspecies interactions have a finite average $\mu/S$ and a variability $\sigma/\sqrt{S}$. The chosen scaling of interaction strengths with $S$ guarantees that the spectral support of the stability matrix [see Eq. (3b) below] does not change with $S$. Formulating the system as in Eq. (1) emphasizes the distribution underlying the random inter-species interactions. It allows us to pinpoint the relevant parameters in the route to chaos.
Fluctuations in interaction strengths between pairs of species are encoded in the components ${\mathbb{A}}_{ij}$ of the interaction matrix $\mathbb{A}$. Being interested in large, heterogeneous ecosystems with no particular structure, we follow a random matrix theory (RMT) approach, May (1972); Allesina and Tang (2015); Akjouj et al. (2024) where the matrix $\mathbb{A}$ belongs to an ensemble of matrices with elements ${\mathbb{A}}_{ij}$ that are normally distributed with vanishing ensemble average, $\langle{\mathbb{A}}_{ij}\rangle=0$, and covariances given by
$\langle{\mathbb{A}}_{ij}{\mathbb{A}}_{kl}\rangle=\delta_{ik}\delta_{jl}+\gamma\delta_{il}\delta_{jk}\,.$ (2)
The cross-diagonal covariance parameter $\gamma\in[-1,1]$ allows to tune the types of two-species interactions that are likely to be encountered in the model. For $\gamma=1$, the interaction matrix $\mathbb{A}$} is totally symmetric and therefore interactions are either mutualistic (${\mathbb{A}}_{ij}\leq 0$ and ${\mathbb{A}}_{ji}\leq 0$), or competitive (${\mathbb{A}}_{ij}\geq 0$ and ${\mathbb{A}}_{ji}\geq 0$). For $\gamma=-1$, on the other hand, $\mathbb{A}$} is totally antisymmetric and all interactions are predator-prey (with ${\mathbb{A}}_{ij}\leq 0$ and ${\mathbb{A}}_{ji}\geq 0$ of vice-versa). Tuning $\gamma$ between $1$ and 1-1, one goes from a system dominated by mutualistic or competitive interaction ($\gamma$ close to $1$) to a system dominated by predator-prey interactions ($\gamma$ close to 1-1). gam Even though the mean $\mu$ introduces a bias towards mutualistic/competitive interactions, its contribution becomes negligible with respect to the contribution of the variability $\sigma$ for large $S$, due to the respective scalings in $S$ of the corresponding terms in Eq. (1). For $\gamma=0$, $\mathbb{A}$ belongs to the Ginibre ensemble or random matrices. Ginibre (1965)
The generalized Lotka-Volterra model of Eq. (1) assumes that $N_{i}(t)$ is real and varies continuously. This is a legitimate assumption only as long as $N_{i}(t)$ is sufficiently large. Large fluctuations have been observed in numerical simulations of Eq. (1), where some species resurrect from minuscule abundances, effectively corresponding to extinction. As a matter of fact, the dynamics of Eq. (1) leads to species extinction only for asymptotically long times, and the standard procedure to solve this atto-fox problem Mollison (1991) is to introduce a small, but finite extinction threshold $N_{c}$. When $N_{i}(t_{c})<N_{c}$, extinction occurs and $N_{i}(t>t_{c})\equiv 0$. We will set $N_{c}=10^{-20}$ but have checked that other choices lead to the same conclusions as presented below.
The model of Eq. (1) has been the focus of many recent investigations in theoretical ecology, yet, it is important to keep in mind the assumptions on which it is based. First, it neglects spatial structures present in real ecological networks. Second, it assumes that interspecies interactions are randomly distributed and neglects higher-order interactions involving three or more species. Third, and finally, a linear intrinsic growth rate $\kappa_{i}N_{i}$ is assumed, whereas some empirical data suggest a sublinear growth rate, which may have a significant impact on stability vs. biodiversity. Hatton et al. (2024); Aguadé-Gorgorió et al. (2025)
From now on, our focus is on the parameters $\sigma$ and $\gamma$ defining the variability of the interspecies interactions and accordingly we fix the growth rates, the initial number of species and the average interaction at values $\kappa_{i}\equiv 1$, $S=157$ and $\mu=5$. We have checked that varying $\mu$ does not change our conclusions as long as $\mu>0$.
Figure 1: Top: abundances for 8 of the $N_{s}=61$ surviving species oscillating in a limit cycle for a realization of the model of Eq. (1), with $\sigma=4.05$ and $\gamma=-0.5$. Bottom: Evolution of the eigenvalues of the stability matrix $\mathbb{M}$} for $\sigma\in[0,4.02]$ (more faded colors indicate smaller values of $\sigma$). Stability of the fixed point is lost when a pair of complex-conjugated eigenvalues cross the imaginary axis, resulting in a Hopf bifurcation and the emergence of the limit cycle shown in the top panel.
The RMT approach to ecosystem dynamics dates back at least to May’s seminal work. May (1972) Fixed-points $\vec{N^{*}}$ of Eq. (1) are defined by $\dot{N}_{i}^{*}=0$ and the dynamics in their vicinity is governed by the stability matrix $\mathbb{M}$
$\delta\dot{\vec{N}}$ $=$ $\mathbb{M}\,\delta\vec{N}\,$, (3a)
$\mathbb{M}_{ij}$ $=$ $-N_{i}^{*}\left(\delta_{ij}+\frac{\mu}{S}+\frac{\sigma}{\sqrt{S}}{\mathbb{A}}_{ij}\right)\,.$ (3b)
[…]
$\sigma$ $<(1+\gamma)^{-1}\,.$ (4)
Beyond that border, fixed-point stability may be recovered via species extinctions, $N_{i}^{*}\rightarrow 0$, except for a number $N_{s}<S$ of surviving species, because then, the stability matrix $\mathbb{A}^{(r)}$ is reduced by removing rows and lines from $\mathbb{A}$, corresponding to the extinct species. Then, the spectrum of $\mathbb{A}^{(r)}$ has an elliptic support with semi-axes $\sigma(1\pm\gamma)\sqrt{N_{s}/S}$. Stability then bounds the number of surviving species as
$N_{s}/S\leq[\sigma(1+\gamma)]^{-2}\,.$ (5)
Note that this argument says nothing about the bifurcation through which fixed-points lose their stability.
Figure 2: Emergence of a limit cycle and transition to chaos for the model of Eq. (1), with $\gamma=-0.5$. The time-evolution of the species abundance $N_{33}$ is plotted for (a) $\sigma=2.5$, (b) $\sigma=2.6$, (c) $\sigma=2.8$, (d) $\sigma=2.85$, and (e) $\sigma=3.19$. Panel (f) shows the numerically computed largest Lyapunov exponent $\lambda$, with green crosses corresponding to the cases shown in panels (a-e). In panel (d) and (e), two initially nearby sets of abundances diverge from one another (black and red curves), reflecting the corresponding positive $\lambda$. Different trajectories repeat similar patterns, albeit in different sequences and without periodicity. Together with $\lambda>0$, this is characteristic of the presence of a strange attractor. There is no such sensitivity to initial conditions in the three other cases, where $\lambda<0$ when the dynamics is attracted to a fixed point [as in panel (a)], and $\lambda=0$ (grey area) in the presence of a limit cycle [as in panels (b) and (c)].
III Results
For $\gamma\neq 1$ the eigenvalues of the real, asymmetric matrix $\mathbb{M}$ are either real, or come in complex-conjugated pairs. This opens up the possibility that the fixed-point loses its stability through a Hopf bifurcation, after which the ecosystem dynamics is attracted to a stable limit cycle. Such a Hopf bifurcation is illustrated in Fig. 1, where the top panel shows periodic oscillations in abundances of the surviving species which are related in the bottom panel to the crossing of the imaginary axis by a complex-conjugated pair of eigenvalues. For large random matrices with $\gamma=0$ it has been shown that only ${\cal O}(\sqrt{S})$ of the $S$ eigenvalues are real, Edelman, Kostlan, and Shub (1994); Edelman (1997); Kanzieper and Akemann (2005) which suggests that such Hopf bifurcations are the rule rather than the exception in ecosystems with high biodiversity.
[…]
The prevalence of limit cycles depends on both the variability $\sigma$ of interspecies interaction and on the off-diagonal covariance parameter $\gamma$. As a matter of fact, to have a limit cycle, one needs a large enough $\sigma$ to have a bifurcation in the first place, moreover, the bifurcation must also involve a pair of complex-conjugated eigenvalues of the stability matrix. Both the value of $\sigma$ and the probability that a complex-conjugated pair of eigenvalues triggers the bifurcation depend on $\gamma$. One expects that more negative values of $\gamma$, together with the associated larger value of $\sigma$ to have a bifurcation favor the occurence of limit cycles. This is confirmed in Fig. 3, which shows the probability $P(\gamma)$ that the ecosystem has reached a state of persistent dynamics for given values of $\sigma$ and $\gamma$. Fig. 3 also confirms that for $\gamma<\sigma^{-1}-1$, below the threshold given by Eq. (4), one is in the regime with a unique asymptotically stable fixed point, without persistent dynamics. The numerical detection method employed for this statistical analysis cannot differentiate between limit cycles and chaotic motion on a strange attractor, because this would require to evaluate the largest Lyapunov exponent in each case, which is computationally time-consuming. However sampling those data indicate that chaotic motion sets in only at larger interactions, and constitutes a significant fraction of the data only for $\sigma=4$. Work to better quantify this fraction is currently underway and goes beyond the scope of this article. Data for larger $\sigma$ are not shown as they exhibit mass extinctions, with only very few surviving species at best, except for the smallest values of $\gamma$.
Figure 3: Probability to find the system in a state of persistent dynamics – either a limit cycle or a state of chaotic dynamics on a strange attractor – as a function of $\gamma$, and for $\sigma=1.5$ [Panel(a)], 2 [(b)], 3 [(c)] and 4 [(d)]. Data have been calculated over ensembles of 100 different interaction matrices, each with 5 different initial conditions. The grey area indicates the range of $\gamma$ where the system converges toward a unique stable fixed point, according to Eq. (4).
The observed increased probability to find limit cycles at negative values of $\gamma$ is related to the associated increased fraction of pairs of complex-conjugated eigenvalues in the spectrum of the stability matrix. Neglecting inhomogeneities in the fixed-point abundances $N_{i}^{*}$, this spectrum is real for $\gamma=1$ and purely imaginary for $\gamma=-1$, and it is expectable that the fraction of real eigenvalues decreases as $\gamma$ decreases. To the best of our knowledge, the only mathematically rigorous result at intermediate values of $\gamma$ is that only a fraction ${\cal O}(\sqrt{S})$ of the eigenvalues are real for the Ginibre ensemble, i.e., at $\gamma=0$Edelman, Kostlan, and Shub (1994); Edelman (1997); Kanzieper and Akemann (2005) While these observations suggest more frequent occurences of limit cycles as $\gamma$ decreases, what truly matters is whether the extreme eigenvalue with largest real part is real or complex, for a fixed realization of the interaction matrix in Eq. (1) with the covariance of Eq. (2). Calculating how likely that is, as a function of $\gamma$, would be a formidable task and we are unaware of any rigorous result in this direction. We therefore resort to numerical calculations (see supplementary material, Fig. S1). As expected, we find that the probability of a complex extreme eigenvalue increases with decreasing $\gamma$. To translate this result into a probability to have a limit cycle emerge through a Hopf bifurcation, we still need to take into account that there is no bifurcation at $\gamma=-1$, since there, increasing $\sigma$ only stretches the spectrum in the imaginary direction. We therefore expect that moderately negative values of $\gamma$ should favor the emergence of limit cycles. This qualitative argument is confirmed by the data shown in Fig. 3. A second numerical result shown in Fig. 5 (see supplementary material) is that the probability that the extreme eigenvalue is complex increases with the number $S$ of species.
Figure 4: Distribution of the number of surviving species for the model of Eq. (1), with $\sigma=2$, $\gamma=0$ (left) and $\sigma=4$, $\gamma=-0.5$ (right). Black histograms correspond to fixed point solutions and red histograms to limit cycle solutions. Distributions are calculated over 400 realizations of the interaction matrix ${\mathbb{A}}_{ij}$, each with 25 different initial abundances. In all considered cases, the system converges to either a fixed point or a limit cycle. Distributions correspond to 22.5% (left panel) and 47.6 % of cycles (right panel).
We translate these results into the language of theoretical ecology. Noting that $\gamma$ interpolates between ecosystems consisting purely of pairs of either mutualistic or competitive species for $\gamma=1$, to ecosystems with only predator-prey pairs of species for $\gamma=-1$ gam , persistent oscillating behavior in population abundances are expected to be prevailing in large, sufficiently interacting ecosystems with a majority of predator-prey pairs. We note that the occurence of Hopf bifurcations is of importance for theoretical ecology. It extends the stability of population coexistence beyond the loss of fixed-point stability, without the need for species extinction. The resulting increasing biodiversity is illustrated in Fig. 4. It is obvious that the average number of coexisting species at fixed $\sigma$ is higher for ecosystems equilibrating to limit cycles ($\langle N_{s}\rangle=58.4$ and $63$) than to fixed-points ($\langle N_{s}\rangle=51.8$ and $53.7$).
IV Conclusions
It has long been known that multi-species ecosystems described by sets of coupled ordinary differential equations of the type given in Eq. (1) may exhibit any dynamical behavior. Smale (1976) Here we have emphasized a route joining the previously observed phases governed by attractive fixed-points to those exhibiting persistent dynamics. Because species in ecosystems and food webs interact with one another in a necessarily asymmetric way, fixed-point instabilities may occur via a Hopf bifurcation. The resulting limit cycles preserve biodiversity, since at least for some range of interaction variability, their stability does not necessitate species extinctions.
Models of theoretical ecology are not expected to precisely reflect ecosystem behaviors. Instead, they may shed qualitative light on different, observed behaviors of real or lab-controlled ecosystems at a statistical level. In that respect, our work emphasizes the existence of various dynamical behaviors, depending on the two parameters $\sigma$ and $\gamma$ governing interspecies interactions. Recent analysis of population time series found evidence of chaotic behavior in at least 30% of the studied populations. Rogers, Johnson, and Munch (2022) Fig. 2 further emphasizes parametrically sizable regimes with small Lyapunov exponents, which has also been observed. Rogers, Johnson, and Munch (2022); Ellner and Turchin (1995) Other regimes with strong sensitivity to even minor parametric changes [e.g., at $\sigma\gtrsim 3$ in Fig. 2(f)] are characteristic of systems close to criticality as discussed in Refs. Solé et al., 2002; Mora and Bialek, 2011. Since criticality slows down the dynamics, the behavior of ecosystems in this latter regime is transient by nature and not governed by any long-time asymptotic. That ecosystems are governed by transient behaviors has been postulated in Refs. Hu et al., 2022; Hastings and Higgins, 1994; Schreiber, 2003; Morozov, Banerjee, and Petrovskii, 2016; Hastings et al., 2018. The emergence of Hopf bifurcations at negative values of $\gamma$ finally explains recent results where ecosystems with dominating predator-prey interactions display oscillatory behaviors in their population dynamics. Mambuca, Cammarota, and Neri (2022)
The present work illustrates that all the different, observed or theoretically postulated behaviors naturally emerge from a single, unified model, without the need for exogeneous intervention. The task at hand now is obviously to try and determine, at least qualitatively, model parameters corresponding to specifically observed ecosystems. Our work has simplified that task in that it identified the two key model parameters $\sigma$ and $\gamma$ driving transitions between different dynamical behaviors.
There are several important direction in which our work should be extended. Among them, we mention first, that we are currently investigating the frequency of occurence of strange attractors vs. restabilization after species extinction at large $\sigma$. Second, further investigations should extend our results to interaction matrices reflecting more realistic topologies of known ecological networks. Fried, Shnerb, and Kessler (2017) Third, investigating species distributions may give precious information on the conditions under which biodiversity and rarity may coexist. Solé et al. (2002) Fourth, ecosystem parameters are modified by climate changes. Rall et al. (2010); Englund et al. (2011); Parain et al. (2019) Investigating changes in ecosystem functioning under climate-induced changes in trophic interactions is of paramount interest. It would evidently have a strong influence on ecosystem functioning in the critical and chaotic regimes with strong parameter sensitivity. There are certainly many other interesting extensions.
Supplementary Material
Section I of the supplementary material provides a numerical estimate of the probability that the extreme eigenvalue of the matrix in Eq. (2) is complex. In Sec. II of the supplementary material, we numerically illustrate various dynamical behaviors that are observed in the system of Eq. (1).
Acknowledgments
We thank Christian Mazza and Xavier Richard for introducing us to theoretical ecology and population dynamics. RD was supported by the Swiss National Science Foundation, under grant nr. 200021_215336.
Author declaration
Conflict of interest
The authors have no conflict to disclose.
Author contributions
Robin Delabays: Conceptualization (support); Formal analysis (support); Investigation (support); Methodology (support); Visualization (equal); Writing – original draft (support); Writing – review & editing (equal).
Philippe Jacquod: Conceptualization (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Visualization (equal); Writing – original draft (lead); Writing – review & editing (equal).
Data availability statement
The data that support the findings of this study are available within the article.
Supplementary material
IV.1 Extreme eigenvalue of a real asymmetric random matrix
The occurence of a Hopf bifurcation requires that the first eigenvalue to cross the imaginary axis has a finite imaginary part – in which case one actually has a pair of different, complex-conjugated eigenvalue. Therefore, the probability of a Hopf bifurcation is larger, if the probability that the extreme eigenvalue of the stability matrix has a finite imaginary part is larger. Fig. 5 shows this probability for the matrix $\mathbb{A}$ of Eqs. (1–2) as a function of $\gamma$ and for different matrix sizes. The probability is larger at more negative values of $\gamma$, moreover, it increases with the system size. Once concludes that Hopf bifurcations are the rule rather than the exception for sufficiently large ecosystems with $\gamma<0$.
Figure 5: Numerically computed probability that the eigenvalue with largest real part of a random matrix defined by Eq. (2) is imaginary, as a function of $\gamma$. Data correspond to averages over exact diagonalizations of 10000 matrices of size $S=57$ (violet), 157 (blue) and 557 (green), and of 5000 matrices with $S=1057$ (red) and 2057 (black). Note the limiting cases (not shown) $P(\gamma=-1)=1$ and $P(\gamma=1)=0$.
IV.2 Long-time asymptotic for different initial conditions
When $\sigma$ is smaller than the May bound, $\sigma<1/(1+\gamma)$, the long-time asymptotic is a single, globally attractive fixed point. We illustrate that, at larger $\sigma$, and with a finite (though small) extinction threshold, $N_{c}=10^{-20}$, different asymptotic behaviors can be reached. Fig. 6 shows five different dynamics obtained from five different initial distribution of populations subjected to the same interaction matrix ${\mathbb{A}}_{ij}$. For clarity, we show only five species for each case. Two initial conditions converge toward a limit cycle and three toward a fixed point. The three fixed points are evidently different. The two cycles differ mostly by the survival of species #75 in panel (a) and its extinction in panel (e). As a consequence, the oscillations are similar, but with larger amplitude in panel (e). We stress that the dynamics has been investigated for much longer times than shown, and that in all cases, the dynamics remain the same as for $t\gtrsim 350$, in particular with the same amplitudes of oscillations in panels (a) and (e).
Setting a higher extinction threshold $N_{c}=10^{-10}$ makes species #75 disappear also in panel (a), however it further modifies the dynamics so that both the cases of panel (a) and (e) converge toward different fixed points (not shown). While the specifically followed dynamics depends on $N_{c}$, the general conclusion that different dynamics can be followed by different initial conditions is general.
Figure 6: Panels (a-e): Five different dynamics, corresponding to the same realization of the interaction matrix, with five different initial distribution of populations. The limit cycles reached in panels (a) and (e) differ by the presence of one additional species in panel (a). Species oscillate about almost the same average values, with however large amplitude in panel (e). Panel (f): Dynamics of species #75 which persists in panel (a) but goes extinct in panel (e). Changing the extinction threshold to $N_{c}=10^{-10}$ leads to two different fixed points for the initial conditions of panels (a) and (e). Five species are shown for each panel. The number of surviving species is $N_{s}=57$ [panel (a)], 51 [panel (b)], 49 [panel (c)], 46 [panel (d)] and 56 [panel (e)].